OpenStudy (anonymous):

use substitute to solve y=2x-2 y=x+2

OpenStudy (solomonzelman):

\[y=2x-2\]\[y=x+2\]\[\color{red} {(x+2)}=2x-2\]\[x+2=2x-2\]\[x+2\color{red} {-x+2}=2x-2\color{red} {-x+2}\]\[4=x\]

OpenStudy (solomonzelman):

\[\color{blue} {MAKES~~~SENSE~~?}\]

OpenStudy (anonymous):

thanks a bunch just wanted to make sure I did it right

OpenStudy (solomonzelman):

Anytime!